# arXiv:2602.14013v2[math.CV] 01 Mar 2026

### v2

@@ 5 Main Lemma for mappings with a branching
Proof of Theorem 1. Let us prove the theorem by contradiction. Assume the assertion of Theorem 1 is false. Then for some $\varepsilon_{0}>0$ and for numbers $\delta_{k}=\frac{1}{k},$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},y_{k}\in D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(y_{k}))\geqslant\varepsilon_{0}$ whenever $|x_{k}-y_{k}|<\frac{1}{k}.$ By Lemma 2 every $f_{m_{k}}$ has a continuous extension to $\partial D_{m_{k}},$ so that $f_{m_{k}}$ is uniformly continuous in $\overline{D_{m_{k}}}.$ Thus, we may consider that the sequence $m_{k},$ $k=1,2,\ldots,$ is increasing. Without loss of generality, we may consider that the above holds for $m=1,2,\ldots,$ but not for $m_{k}.$ In particular,
Now, by the triangle inequality,
By the compactness of $\overline{{\mathbb{R}}^{n}},$ we may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ There are two cases: 1) $x_{0}\in D_{0},$ 2) $x_{0}\in\partial D_{0}.$ In the first case, when $x_{0}$ is the inner point of $D_{0},$ all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [Sev2, Lemma 4.2] and Proposition 2 for $p=n$ and [Sev2, Theorems 4.1, 4.3] for n1<p<n.n-1<p<n. In this case, the right side in (4.2) tends to zero as $m\rightarrow\infty$ which contradicts with (4.1), and the proof finishes.
Let us consider the case 2) $x_{0}\in\partial D_{0}.$ Observe that $A_{m}:=A\subset{\mathbb{R}}^{n}\setminus B(x_{0},\varepsilon_{0})$ for some $\varepsilon_{0}>0,$ because the $A$ is a continuum in $D_{0}.$ Thus, all of conditions of Theorem 2 are satisfied. In particular, it follows from this theorem that $f_{m}$ has a continuous extension to $\partial D_{m}$ for any $m\in{\mathbb{N}}.$ Thus, there is $x_{m}\in D_{m}$ such that $h(f_{m}(x^{\,\prime}_{m}),f_{m}(x_{0}))<\frac{1}{m}.$ Now, by the triangle inequality, by (4.2) and by Theorem 2 we obtain that
$h(f_{m}(x_{m}),f_{m}(y_{m}))\leqslant h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(x_{0}))+$
$+h(f_{m}(x_{0}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(y_{m}))<$ (5.7)
$+h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))+h(f_{m}(x^{\,\prime}_{m}),f_{m}(y_{m}))+\frac{2}{m}\rightarrow 0\qquad\text{as}\qquad m\rightarrow\infty.$
The relations (5.6) and (5.7) contradict each other and prove the uniformly equicontinuity of $f_{m}.$
[…]
@@ 6 The versions for prime ends
The following statement may be proved completely similarly to [Sev2, Lemma 5.2]. For the sake of consistency in presentation, we present its proof in full.
Lemma 6.1.  Let $p\geqslant 1,$ let $D$ be a quasiconformally regular domain, let $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete and closed ring $Q$-mapping with respect to the $p$-modulus at all points $b\in\partial D,$ let $f(D)=D^{\,\prime}$, and let $D^{\,\prime}$ has a strongly accessible boundary with a respect to $p$-modulus at least at one point $y\in C(f,b).$ Assume that, for any $b\in\partial D$ there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$ (6.1)
@@
$\int\limits_{A(\varepsilon,\varepsilon_{0},b)}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$, (6.2)
@@
Proof.   Put $P\in E_{D}.$ Since $\overline{D^{\,\prime}}$ is a compactum in $\overline{{\mathbb{R}}^{n}},$ we need to prove that the cluster set
$L=C(f,P):=\Big\{y\in{{\mathbb{R}}^{n}}\colon y=\lim_{k\rightarrow\infty}f(x_{k}),x_{k}\rightarrow P,x_{k}\in D\Big\}$
consists of a unique point $y_{0}\in\partial D^{\,\prime}$. Observe that $L\neq\varnothing$ because $\overline{D^{\,\prime}}$ is compact, and $L$ is a subset of $\partial D^{\,\prime}$ because $f$ is closed (see [Vu, Theorem 3.3]). Suppose that there are at least two points $y_{0}$ and $z_{0}$ in $L$. That is, there are at least two sequences $x_{k}$, $x^{\,\prime}_{k}\in D$ with $x_{k}\rightarrow P$ and $x^{\,\prime}_{k}\rightarrow P$ as $k\rightarrow\infty$ and $f(x_{k})\rightarrow y_{0}$ and $f(x^{\,\prime}_{k})\rightarrow z_{0}$ as $k\rightarrow\infty$. By [KR, Lemma 2], every prime end $P$ of a regular domain $D$ in ${\mathbb{R}}^{n}$, $n\geqslant 2$, contains a chain of cuts $\sigma_{k}$ that lie on the spheres $S_{k}$ centered at some $x_{0}\in\partial D$ and whose Euclidean radii $r_{k}$ tend to zero as $k\rightarrow\infty$. Let $D_{k}$ be the domain associated with $\sigma_{k}$, $k=1,2,\dots$. Without loss of generality, passing if necessary to a subsequence, we may assume that $x_{k}$, $x^{\,\prime}_{k}\in D_{k}$. Indeed, since the sequences $x_{k}$ and $x^{\,\prime}_{k}$ converge to the prime end $P$, there is a number $k_{1}\in{\mathbb{N}}$ with $x_{k_{1}}$, $x^{\,\prime}_{k_{1}}\in D_{1}$. Next, there exists $k_{2}\in{\mathbb{N}}$, $k_{2}>k_{1}$, with $x_{k_{2}}$, $x^{\,\prime}_{k_{2}}\in D_{2}$, and so on. In general, at the $m$th step we find $k_{m}\in{\mathbb{N}}$, $k_{m}>k_{m-1}$ with $x_{k_{m}}$, $x^{\,\prime}_{k_{m}}\in D_{m}$. Continuing in this way, we obtain two sequences $x_{k_{m}}$ and $x^{\,\prime}_{k_{m}}$ whose $m$th terms belong to $D_{m}$ and that converge to $P$ as $m\rightarrow\infty$ and, moreover, $f(x_{k_{m}})\rightarrow y_{0}$ and $f(x^{\,\prime}_{k_{m}})\rightarrow z_{0}$ as $m\rightarrow\infty$. Redenoting, if necessary, $x_{k_{m}}$ by $x_{m}$, we obtain a sequence $x_{m}$ with the required properties.
We note that $y_{0}$, $z_{0}\in\partial D^{\,\prime}$ because, by assumption, $C(f,\partial D)\subset\partial D^{\,\prime}$. By the definition of a strongly accessible boundary relative to the $p$-modulus at $y_{0}\in\partial D^{\,\prime}$, for every neighborhood $U$ of this point there exists a compact set $C_{0}^{\,\prime}\subset D^{\,\prime}$, a neighborhood $V$ of $y_{0}$ with $V\subset U$, and a number $\delta>0$ such that
$M_{p}(\Gamma(C_{0}^{\,\prime},F,D^{\,\prime}))\geqslant\delta>0$ (6.3)
for every continuum $F$ intersecting $\partial U$ and $\partial V$. The assumption $C(f,\partial D)\subset\partial D^{\,\prime}$ shows that for $C_{0}:=f^{\,-1}(C_{0}^{\,\prime})$ we have $C_{0}\cap\partial D=\varnothing$ (see [Vu, Theorem 3.3]). We may consider that $d(x_{0},C_{0})\geqslant\varepsilon_{0}.$ Since $I(P)=\bigcap_{m=1}^{\infty}\overline{D_{m}}\subset\partial D$ (see [KR, Proposition 1]), without loss of generality we may assume that $C_{0}\cap\overline{D_{k}}=\varnothing$ for every $k\in{\mathbb{N}}$. We join $x_{k}$ and $x^{\,\prime}_{k}$ by a paths $\gamma_{k}$ within $D_{k}$. Observe that $f(x_{k})\in V$ and $f(x^{\,\prime}_{k})\in D\setminus\overline{U}$ for all sufficiently large $k\in{\mathbb{N}}$. Thus, there exists $k_{0}\in{\mathbb{N}}$ such that, by (6.3),
$M_{p}(\Gamma(C^{\,\prime}_{0},|f(\gamma_{k})|,D^{\,\prime}))\geqslant\delta>0$ (6.4)
for all $k\geqslant k_{0}\in{\mathbb{N}}$. We denote by $\Gamma_{k}$ the family of all half-open paths $\beta\colon[a,b)\rightarrow{\mathbb{R}}^{n}$ such that $\beta(a)\in|f(\gamma_{k})|$, $\beta(t)\in D^{\,\prime}$ for all $t\in[a,b)$ and, moreover, $\lim\limits_{t\rightarrow b-0}\beta(t):=B\in C^{\,\prime}_{0}$. Clearly,
$M_{p}(\Gamma_{k})=M_{p}(\Gamma(C^{\,\prime}_{0},|f(\gamma_{k})|,D^{\,\prime})).$ (6.5)
For every fixed $k\in{\mathbb{N}}$ with $k\geqslant k_{0}$, we consider the family $\Gamma^{\,\prime}_{k}$ of (full) liftings $\alpha\colon[a,b)\rightarrow D$ of the family $\Gamma_{k}$ with origin in $|\gamma_{k}|$, i.e., $f\circ\alpha=\beta$, $\beta\in\Gamma_{k}$, and $\alpha(a)\in|\gamma_{k}|$. Since the mapping $f$ is closed, it preserves the boundary (see [Vu, Theorem 3.3]) and, therefore, the path $\alpha$ lies in $D$ together with its closure. Moreover, since $f$ is discrete, $\alpha$ can be extended up to a continuous path $\overline{\alpha}\colon[a,b]\rightarrow D$. Note that $\overline{\beta}(b)\in C^{\,\prime}_{0}$ by definition, so $\overline{\alpha}(1)\in C_{0}$ by the definition of $C_{0}$. Consequently, $\overline{\alpha}\in\Gamma(|\gamma_{k}|,C_{0},D)$, where $\overline{\alpha}$ denotes the extended path $\overline{\alpha}\colon[a,b]\rightarrow D$. We observe that, $\Gamma(|\gamma_{k}|,C_{0},D)>\Gamma(\sigma_{k},S(x_{0},\varepsilon_{0}),D),$ where $\sigma_{k}$ is the cut corresponding to $D_{k}$. Thus, the definition of the ring $Q$-mapping (1.3) may be applied to the family of paths $\Gamma(\sigma_{k},S(x_{0},\varepsilon_{0}),D).$ In this connection, we observe that $\sigma_{k}\subset S(x_{0},r_{k})$ for some $x_{0}\in\partial D$ and some sequence $r_{k}>0$, $r_{k}\rightarrow 0$ as $k\rightarrow\infty$ (see [KR, Lemma 2]). Next, we observe that the function
$\eta_{k}(t)=\begin{cases}\psi(t)/I(r_{k},\varepsilon_{0}),&t\in(r_{k},\varepsilon_{0}),\\ 0,&t\in{\mathbb{R}}\setminus(r_{k},\varepsilon_{0}),\end{cases}$
where
$I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt,$
satisfies a normalization condition of the form (1.4). By what has been proved, we have
$\Gamma^{\,\prime}_{k}\subset\Gamma(|\gamma_{k}|,C_{0},D),$
whence $M_{p}(f(\Gamma_{k}^{\,\prime}))\leqslant M_{p}(f(\Gamma(|\gamma_{k}|,C_{0},D)))$. Therefore, by the definition of a ring $Q$-mapping at a boundary point relative to the $p$-modulus, and also by (1.3)–(1.4), we have
$M_{p}(f(\Gamma^{\,\prime}_{k}))\leqslant M_{p}(f(\Gamma(|\gamma_{k}|,C_{0},D)))\leqslant M_{p}(f(\Gamma(\sigma_{k},S(x_{0},\varepsilon_{0}),D))\leqslant\Delta(k),$ (6.6)
where $\Delta(k)\rightarrow 0$ as $k\rightarrow\infty$. However, $\Gamma_{k}=f(\Gamma^{\,\prime}_{k})$, and by (6.6) we obtain
$M_{p}(\Gamma_{k})=M_{p}\left(f(\Gamma^{\,\prime}_{k})\right)\leqslant\Delta(k)\rightarrow 0$ (6.7)
as $k\rightarrow\infty$. But (6.5) and (6.7) contradict (6.4), which proves the lemma.  $\Box$  $\Box$
The following statement holds.
@@
Proof.   We will mainly use the methodology that has been used for mappings of a fixed domain, see, e.g., [SevSkv]. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in D_{m_{k}}$ such that $\rho(x_{k},x^{\,\prime}_{k})<1/k,$ however, $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ By the definition, $\overline{D_{0}}_{P}$ is a compact space, so that we may consider that the sequences $x_{k},x^{\,\prime}_{k}\rightarrow P_{0}\in\overline{D_{0}}_{P}$ by the metrics $\rho.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that $P_{0}\in\overline{D_{k_{0}}}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow P_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 6. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e., $x_{m},x^{\,\prime}_{m}\in D_{m}$ and
@@
If $P_{0}\in D_{0},$ then all mappings $f_{m}$ except of finite numbers are defined in some ball $B(x_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $x_{0}$ (see [RS, Lemma 4.9] for $p=n$ and [Sev2, Lemma 4.3] for n1<p<n.n-1<p<n. Now, we may consider that $P_{0}\in E_{D_{0}}=\overline{D_{0}}_{P}\setminus D_{0}.$
Let $\sigma_{m},$ $m=0,1,2,\ldots,$ lie on the spheres $S(z_{0},r_{m})$ centered at some point $z_{0}\in\partial D_{0},$ where $r_{m}\rightarrow 0$ as $m\rightarrow\infty$ (such a sequence $\sigma_{m}$ exists by [IS, Lemma 3.1], cf. [KR]). Since $D_{m}$ is a regular sequence of domains with a respect to $D_{0}$ in the terms of prime ends, there exists a sequence of cuts $\varsigma_{m}$ and domains $d^{\,\prime}_{m},$ equivalent to $\sigma_{m}$ and $d_{m},$ respectively, such that the following condition holds: given $d^{\,\prime}_{k},$ $k=1,2,\ldots,$ there is $M_{1}=M_{1}(k)$ such that $d^{\,\prime}_{k}\cap D_{m}$ is a non-empty connected set for every $m\geqslant M_{1}(k).$ Now, we will consider that $d^{\prime}_{k}\subset d_{k}$ for any $k=1,2,\ldots.$
@@
for every Lebesgue measurable function $\eta:(\frac{1}{2^{m}},\varepsilon_{0})\rightarrow[0,\infty]$ such that $\int\limits_{r_{m}}^{r_{1}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
@@
where $I(\varepsilon,\varepsilon_{0}):=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt.$ Therefore, it follows from (2.4) and (6.11) that
@@
The existing of a subsequence $f_{m_{k}},$ $k=1,2,\dots,$ which converges to $f$ locally uniformly in $D_{0}$ with respect to $\rho,$ and the continuity of $f$ in $\overline{D_{0}}_{P}$ may be proved similarly to how it was done under the proof of Theorem 1. The relation $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots$ and some $\delta_{1}>0$ follows by Lemma 3.
Finally, let $A_{0}$ be some another compactum in $D_{0}.$ Now, by Corollary 3 there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ for all $m\geqslant M_{0}$ and $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ The proof is complete. $\Box$
The obtained contradiction completes the proof. $\Box$
@@
Proof.   The proof is much the same as that of Lemma 6, so we will limit ourselves to a schematic proof.

### v3

@@ 5 Main Lemma for mappings with a branching
Proof of Theorem 1. The proofs of the statements in item I are based on Theorem 5 (instead of Theorem 2) and are completely similar to the corresponding proofs of Theorem 1.
II. Let $A_{0}$ be a compactum in $D_{0}.$ We show that there is $\delta_{2}>0$ and $M_{0}\in{\mathbb{N}}$ such that $A_{0}\subset D_{m}$ and $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0}.$ Note that, the relation $A_{0}\subset D_{m}$ for sufficiently large $m\in{\mathbb{N}}$ holds obviously, because $D_{m}$ converges to $D_{0}$ as its kernel by the assumption. Now, we show that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for sufficiently large $m\in{\mathbb{N}}.$ Assume the contrary. Then $h(f_{m_{k}}(A_{0}),\partial f_{m_{k}}(D_{m_{k}}))\rightarrow 0$ as $k\rightarrow\infty$ for some increasing a sequence of numbers $m_{k},$ $k=1,2,\ldots.$ Going to renumbering, if required, we may consider that the latter holds for $m$ instead of $m_{k},$ i.e.,
$h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\rightarrow 0\,,\quad m\rightarrow\infty\,.$ (5.5)
Let $K_{m}\subset D^{\,\prime}_{m}$ be a continuum from the definition of the class $\mathfrak{R}_{Q,\delta,p,E}(D_{0},\mathfrak{D})$ such that $h(K_{m})\geqslant\delta$ and $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0.$ Let $h(K_{m})=h(y_{m},z_{m}),$ where $y_{m},z_{m}\in K_{m}.$ Now, $y_{m}=f_{m}(x_{m})$ and $z_{m}=f_{m}(x^{\,\prime}_{m})$ for some $x_{m},x^{\,\prime}_{m}\in f_{m}^{\,-1}(K_{m}).$ We may consider that $x_{m}\rightarrow x_{0}$ and $x^{\,\prime}_{m}\rightarrow x^{\,\prime}_{0}$ as $m\rightarrow\infty.$ Due to the condition $h(f_{m}^{\,-1}(K_{m}),\partial D_{m})\geqslant\delta>0$ we have that $h(x_{m},\partial D_{m})\geqslant\delta$ and $h(x^{\,\prime}_{m},\partial D_{m})\geqslant\delta,$ $m=1,2,\ldots.$ Since by the assumption $D_{m}\subset D_{0},$ $m=1,2,\ldots,$ by Proposition 4 $h(x_{m},\partial D_{0})\geqslant\delta$ and $h(x^{\,\prime}_{m},\partial D_{0})\geqslant\delta,$ $m=1,2,\ldots.$ Now, $x_{0},x^{\,\prime}_{0}\in D_{0}.$ Let $\sigma_{1},\sigma_{2}>0$ be numbers such that $\overline{B(x_{0},\sigma_{1})}\subset D_{0}$ and $\overline{B(x^{\,\prime}_{0},\sigma_{2})}\subset D_{0}.$ We may consider that $x_{m}\in\overline{B(x_{0},\sigma_{1})}$ and $x^{\,\prime}_{m}\in\overline{B(x^{\,\prime}_{0},\sigma_{2})}$ for any $m\in{\mathbb{N}}.$ Now, we join the points $x_{0}$ and $x^{\,\prime}_{0}$ by a path $\gamma$ in $D_{0}.$ Let $A:=\overline{B(x_{0},\sigma_{1})}\cup|\gamma|\cup\overline{B(x^{\,\prime}_{0},\sigma_{2})},$ now $A$ is a continuum in $D_{0}.$ Since $D_{m}$ converges to $D_{0}$ as its kernel, we may consider that $A\subset D_{m}$ for any $m=1,2,\ldots.$ Observe that, $h(f_{m}(A))\geqslant h(f_{m}(x_{m}),f_{m}(x^{\,\prime}_{m}))=h(y_{m},z_{m})\geqslant\delta>0.$ Now, by Lemma 3 $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{*}>0$ for all $m\in{\mathbb{N}}$ and some $\delta_{*}>0.$ Besides that, by Corollary 3 $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{**}>0$ for all $m\in{\mathbb{N}}$ and some $\delta_{**}>0.$ The latter contradicts with the assumption (5.5). Thus, $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for sufficiently large $m\in{\mathbb{N}},$ as required.
Let us to prove that, the limit mapping $f$ is boundary preserving. In other words, wee need to prove that, if $x_{0}\in\partial D_{0},$ then $f(x_{0})\in\partial f(D_{0}).$ Assume the contrary: let $x_{0}\in\partial D_{0},$ however, $f(x_{0})\in{\rm Int\,}f(D_{0}).$ Now, $\partial D_{m}\cap U\neq\varnothing$ for any neighborhood $U$ of $x_{0}$ and for infinitely many $m=1,2,\ldots.$ So, we may construct a sequence $x_{k}\in\partial D_{m_{k}},$ $k=1,2,\ldots,$ such that $x_{k}\rightarrow x_{0}$ as $k\rightarrow\infty.$ For simplicity, we may consider that $x_{m}\in\partial D_{m}$ and $x_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Since $f_{m}$ are closed, they are boundary preserving (see, e.g., [Vu, Theorem 3.3]). Thus, for any $m\in{\mathbb{N}}$ we may find $y_{m}\in D_{m}$ such that $|x_{m}-y_{m}|<\frac{1}{m}$ and $h(f_{m}(y_{m}),\partial f_{m}(D_{m}))<\frac{1}{m}.$ Let $h(f_{m}(y_{m}),\partial f_{m}(D_{m}))=h(f_{m}(y_{m}),z_{m}).$ Let $A_{0}:=\{y_{0}\}$ and $y_{0}$ is some point of $D_{0}$ with $f(y_{0})=f(x_{0}).$ Now, $y_{0}\in D_{m}$ for all $m\in{\mathbb{N}}$ except a finite number. By the triangle inequality and proving above we obtain that
$h(f_{m}(y_{0}),z_{m})\leqslant h(f_{m}(y_{0}),f(x_{0}))+h(f(x_{0}),f_{m}(y_{m}))+h(f_{m}(y_{m}),z_{m})\rightarrow 0$
as $m\rightarrow\infty.$ The latter contradicts with the statement that $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0},$
Let us to prove the relation (1.6). We lead the proof by the contradiction, i.e., assume that (1.6) does not hold. Now, given $k\in{\mathbb{N}}$ there is $m=m_{k}\in{\mathbb{N}}$ such that $B_{h}(f_{m_{k}}(x_{0}),1/k)\setminus f_{m_{k}}(D_{m_{k}})\neq\varnothing$ whenever $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}\cap\bigcap\limits_{m=1}^{\infty}D_{m}.$ Let $p_{k}\in B_{h}(f_{m_{k}}(x_{0}),1/k)\setminus f_{m_{k}}(D_{m_{k}}).$ We join the points $f_{m_{k}}(x_{0})$ and $p_{k}$ inside the ball $B_{h}(f_{m_{k}}(x_{0}),1/k)$ by a path $\gamma_{k}.$ Now, $|\gamma_{k}|\cap f_{m_{k}}(D_{m_{k}})\neq\varnothing\neq|\gamma_{k}|\setminus f_{m_{k}}(D_{m_{k}})$ and by Proposition 3 $|\gamma_{k}|\cap\partial f_{m_{k}}(D_{m_{k}})\neq\varnothing.$ Thus, $h(f_{m_{k}}(x_{0}),\partial f_{m_{k}}(D_{m_{k}}))\rightarrow 0$ as $k\rightarrow\infty,$ that contradicts with $h(f_{m}(A_{0}),\partial f_{m}(D_{m}))\geqslant\delta_{2}$ for every $m\geqslant M_{0},$ proved above. Theorem is proved. $\Box$
@@ 6 The versions for prime ends
The following statement may be proved completely similarly to [Sev1, Lemma 5.2], cf. [Ad, Theorem 5, Example 5], [KR, Lemma 5, Theorem 3].
Lemma 6.1.  Let $p\geqslant 1,$ let $D$ be a quasiconformally regular domain, let $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete and closed ring $Q$-mapping with respect to the $p$-modulus at all points $b\in\partial D,$ let $f(D)=D^{\,\prime}$, and let $D^{\,\prime}$ be a domain with a strongly accessible boundary with a respect to $p$-modulus at least at one point $y\in C(f,b).$ Assume that, for any $b\in\partial D$ there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$
@@
$\int\limits_{A(b,\varepsilon,\varepsilon_{0})}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$,
@@
The following statement holds, cf. [SevSkv].
@@
Proof.   I. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in D_{m_{k}}$ such that $\rho(x_{k},x^{\,\prime}_{k})<1/k,$ however, $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ By the definition, $\overline{D_{0}}_{P}$ is a compact space, so that we may consider that the sequences $x_{k},x^{\,\prime}_{k}\rightarrow P_{0}\in\overline{D_{0}}_{P}$ by the metrics $\rho.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that, $P_{0}\in\overline{D_{k_{0}}}_{P}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow P_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 6. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that the relation $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e., $x_{m},x^{\,\prime}_{m}\in D_{m}$ and
@@
If $P_{0}\in D_{0},$ then all mappings $f_{m}$ except of finite numbers are defined in some ball $B(P_{0},\varepsilon_{0})\subset D_{0}\cap D_{m},$ $m\geqslant M_{0}.$ Now, $f_{m}$ is equicontinuous at $P_{0}$ (see [RS1, Lemma 4.9] for $p=n$ and [Sev1, Lemma 4.3] for n1<p<n.n-1<p<n. Therefore, we may consider that $P_{0}\in E_{D_{0}}=\overline{D_{0}}_{P}\setminus D_{0}.$
Let $d_{m}$ and $\sigma_{m},$ $m=0,1,2,\ldots,$ be sequences of domains and cuts in $P_{0},$ respectively, while $\sigma_{m}$ lie on the spheres $S(z_{0},r_{m})$ centered at some point $z_{0}\in\partial D_{0},$ where $r_{m}\rightarrow 0$ as $m\rightarrow\infty$ (such a sequence $\sigma_{m}$ exists by [IS, Lemma 3.1], cf. [KR]). Since $D_{m}$ is a regular sequence of domains with a respect to $D_{0}$ in the terms of prime ends, there exists a sequence of cuts $\varsigma_{m}$ and domains $d^{\,\prime}_{m},$ equivalent to $\sigma_{m}$ and $d_{m},$ respectively, such that the following condition holds: given $d^{\,\prime}_{k},$ $k=1,2,\ldots,$ there is $M_{1}=M_{1}(k)$ such that $d^{\,\prime}_{k}\cap D_{m}$ is a non-empty connected set for every $m\geqslant M_{1}(k).$ Now, we will consider that $d^{\,\prime}_{k}\subset d_{k}$ for any $k=1,2,\ldots.$
@@
for every Lebesgue measurable function $\eta:(r_{m},r_{1})\rightarrow[0,\infty]$ such that $\int\limits_{r_{m}}^{r_{1}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
@@
where $I(r_{m},r_{1}):=\int\limits_{r_{m}}^{r_{1}}\psi(t)\,dt.$ Therefore, it follows from (2.2) and (6.4) that
@@
[…]
@@
Lemma 6.3.   The statement of Theorem 1 remains true, if, under assumptions of this theorem, instead of the assumption 4) mentioned here we require the following: there exists $\varepsilon_{0}=\varepsilon_{0}(x_{0})>0$ and a Lebesgue measurable function $\psi:(0,\varepsilon_{0})\rightarrow[0,\infty]$ such that (2.1) holds while the relation (2.2) is satisfied as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
@@
Proof.   I. The proof is much the same as that of Lemma 6, so we will limit ourselves to a schematic proof.